Total points in a game was 101, 45 goals, how many 2 pointers and how many 3 pointers. I can't seem to figure out the correct equation.
Answer by edjones(8007)

Let x=number of 2 pointers and 45-x=number of 3 pointers.
2x+3(45-x)=101
2x+135-3x=101
-x=-34
x=34
.
Ed